OpenStudy (anonymous):

z2 + 5z – 36 = (z – 4)(z + )

OpenStudy (anonymous):

what you need help with ?

OpenStudy (anonymous):

after the last z its a blank and its asking for a number i guess to make it equivalent

OpenStudy (anonymous):

It will be 9 in the blank

OpenStudy (anonymous):

positive? haha and how did you figure that out?

OpenStudy (anonymous):

z2 + 5z – 36 We will break the middle term z^2+9z-4z-36 z(z+9)-4(z+9) (x-4)(z+9) See this you'll understand :)

OpenStudy (anonymous):

Yes positive 9 :D

OpenStudy (anonymous):

z2 + 4z – 45 = (z – 5)(z + ) what about this one?

OpenStudy (anonymous):

Do you have how to break the middle term?

OpenStudy (anonymous):

no i don't i'm still confused sorry.

OpenStudy (anonymous):

LOL okay I'll explain you :)

OpenStudy (anonymous):

See to break the middle term we should 2 numbers. When we add those numbers it becomes the middle term and when we multiply them it becomes the last term for example x^2+8x+7=0 Here we will break the middle term 1 and 7 will be the numbers x^2+1x+7x+7=0 So if we add 7+1 it becomes 8 i mean the middle term and when we multiply them it becomes the last term that is 7 Got it what I explained? :) If no then just lemme know I'll explain it again :)

OpenStudy (anonymous):

um okay but where do you get the 1?

OpenStudy (anonymous):

z2 + 5z – 36 See this question here I will break the middle term by numbers that are -4 and 9 so when we add them they become 5 and when I multiply them they become -36 Got it?

OpenStudy (anonymous):

no because i dont understand how to get the numbers like -4 and 9?

OpenStudy (anonymous):

Ohh listen see need to do this by our own mind :/

OpenStudy (anonymous):

do you just find two numbers that when you multiply them you get the last number?

OpenStudy (anonymous):

mid term breaking is a part of technique used in factorization of [usually] algebraic expressions & equations. it is also used in solving for quadratic equations. for example in factorization of x^2 -7x +12=0 middle term is splitted to get some common factor. herein x^2 -3x -4x +12=0 --------- x ( x -3) -4(x-3)=0 (x-3) (x-4)=0 so that x = 3 & x=4

OpenStudy (anonymous):

See this. Understood?

OpenStudy (anonymous):

im sorry but no its fine though im just dumb haha but thank you so much for all of your help

OpenStudy (anonymous):

LOL no it was even difficult for me but now I'm every good at this :D

OpenStudy (anonymous):

Easiest way is you take L.c.m of the last term :/ Try I'm sure this will help you :)

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

z2 + 4z – 45 = (z – 5)(z +) so for this one i should use 9 and 5?

OpenStudy (anonymous):

or -5 and 9

OpenStudy (anonymous):

Yes you are right :) You aren't dumb but a little slow but don't worry :)